A sum of money amountaed to RS5940 at 4% per annum in 5 year.find original sum?

Grade:7

1 Answers

Shreya singh
6 years ago
if you want to take out how much money I would be getting with the simple intrest then 
SI= (p*r* t)/ 100
So, SI = (940* 4* 5)/ 100
SI= 118
Total sum = SI + p
Total sum= 118+940
=1128